# Algebraic Limits of Sandpiles

### v2

@@ Abstract
The paper contributes to building algebraic foundations of self-organized criticality answering a previously unsolved question about the limiting structure of the extended sandpile group as well as relating it to another limit at the level of classical sandpile groups with respect to certain monomorphisms, and puts forward a concept of canonical sandpile epimorphisms, drawing an unexpected consequence about the divisibility properties of the numbers of spanning trees on rectangles.
@@ 2 Sandpile groups and harmonic functions
where $/\mathbb{Z}$ takes the fractional part of the extended recurrent state and $\mathbb{R}/\mathbb{Z}$ is the quotient of $\mathbb{R}$ by the additive subgroup $\mathbb{Z}$ which topologically is a circle.
@@
$0\rightarrow\mathcal{H}_{\mathbb{Z}}(\Gamma)\rightarrow\mathcal{H}_{\mathbb{R}}(\Gamma)\rightarrow\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\Gamma)\rightarrow\operatorname{Coker}\Delta^{\circ}\rightarrow 0$
@@
$0\rightarrow\mathcal{H}_{\mathbb{Z}}(\Gamma)\rightarrow\mathbb{Z}^{\partial\Gamma}\rightarrow G(\Gamma)\rightarrow\operatorname{Coker}\Delta^{\circ}\rightarrow 0$
@@
One can approach this by observing first that the restriction homomorphisms to the domain from its diamond hull (the minimal diamond, not necessarily Aztec, containing this domain) is an isomorphism. The surjectivity for two diamonds, where one is inside the other, is obvious. ∎
@@
For $\Gamma\subset\mathbb{Z}^{2}$ the sandpile group $G(\Gamma)$ is generated by adding sand at the boundary.
@@
Note that for any Abelian group $A$
$\mathcal{H}_{A}(\mathbb{Z}^{2})\cong\lim_{\mathbb{Z}^{2}\leftarrow\Gamma}\mathcal{H}_{A}(\Gamma).$
Lemmas 3 and 1 imply exactness of columns, and Lemmas 2 and 4 give exactness of rows in the diagram below, now we can apply (for the fifths and last time) the snake lemma:
00{0}00{0}00{0}${\mathcal{H}_{\mathbb{Z}}(\mathbb{Z}^{2})}$${\prod\mathcal{H}_{\mathbb{Z}}(\Gamma_{n})}$${\prod\mathcal{H}_{\mathbb{Z}}(\Gamma_{n})}$00{0}${\mathcal{H}_{\mathbb{R}}(\mathbb{Z}^{2})}$${\prod\mathcal{H}_{\mathbb{R}}(\Gamma_{n})}$${\prod\mathcal{H}_{\mathbb{R}}(\Gamma_{n})}$00{0}${\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\mathbb{Z}^{2})}$${\prod\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\Gamma_{n})}$${\prod\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\Gamma_{n})}$00{0}00{0}00{0}
@@ 3 Monomorphisms and epimorphisms
[…]
@@
of the epimorphism and the monomorphism is the identity homomorphism multiplied by $mn$. There are two extreme cases: when $mn$ is coprime with $\gamma(p,q)=|G(\Gamma(p,q))|,$ in which case $G(\Gamma(p,q))$ is realized as a direct summand of $G(\Gamma(mp,nq)),$ or when $mn$ is divisible by $\gamma(m,n),$ and therefore we realize the group $G(\Gamma(p,q))$ as a subgroup of the kernel of the epimorphism, which gives the second assertion of the following theorem.
@@
Table 1: Orders of sandpile groups on $(0,m)\times(0,n)\cap\mathbb{Z}^{2}.$ Observe the divisibility properties in accordance with Theorem 2. For instance, $\gamma(2,4)$ is divisible by $\gamma(2,2),$ but not by $(\gamma(2,2))^{2}$ since the scaling factor $2=1\times 2$ is not 00 modulo $\gamma(2,2)=4;$ on the other hand $(\gamma(2,2))^{2}$ divides $\gamma(4,4).$
@@
Let $\gamma(p,q)=|G(\Gamma(p,q))|$ be the number of spanning trees on the rectangular graph with non-sink vertices $\Gamma(p,q).$ Then for any quadruple of positive integers $n,m,p,q$ the number $\gamma(p,q)$ divides the number $\gamma(mp,nq).$ Moreover, if $mn$ is divisible by $\gamma(p,q)$, then $(\gamma(p,q))^{2}$ divides $\gamma(mp,nq).$
@@
There is the following commutative diagram:
${G(\Gamma(m,n))}$${G(\Gamma(pm,qn))}$${\widetilde{G}(\Gamma(m,n))}$${\widetilde{G}(\Gamma(pm,qn)).}$
@@
It is instructive to see what is the situtation in the one-dimensional case, which is as a general rule dramatically simpler. The extended sandpile group of an integer interval $(0,n)\cap\mathbb{Z}$ for $n\geq 3$ is a two-dimensional torus made of circle-valued arithmetic progressions, and the usual sandpile group is a cyclic group of order $n,$ corresponding to such progressions which take $0\in\mathbb{R}/\mathbb{Z}$ at 00 and $n.$ The direct limit of this groups is $\mathbb{Q}/\mathbb{Z},$ a rational circle corresponding to periodic circle-valued arithmetic progressions with value 00 at the origin, and the inverse limit under the epimorphisms is ^$\widehat{\mathbb{Z}}=\prod_{l}\mathbb{Z}_{l},$ a direct product of all additive groups of $l$-adic integers, which indeed has no torsion.
@@ 4 Discussion
There is a striking and not fully explored relation [8] between the sandpile model on a rectangular domain and domino tilings on it. The group playing the role in this relation is not the whole sandpile group of the rectangle, but rather a subgroup of elements symmetric with respect to horizontal and vertical reflections. It seems that with a bit of extra work our Theorem 2 may be adapted to the case of such subgroups, and therefore give the analogous divisibility result for the numbers of domino tilings.
The category $\mathcal{D}\mkern-12.0mu\mathcal{D}$ which was passingly introduced in Section 2 imitates by its essence a category of open subsets of some topological space. It makes sense to extend its class of morphisms, leaving the objects intact, by allowing all maps between graphs whose pullbacks on integer-valued functions commute with $\Delta^{\circ}.$ Morally, this allows for coverings, thus we may denote such a category as $\mathcal{D}\mkern-12.0mu\mathcal{D}^{et}$ with $et$ standing for “étale”, and $\mathcal{H}_{A}$ is still some kind of a sheaf. It is conceivable that it would be fruitful to study “the universal extended sandpile group” defined as the inverse limit (i.e. global sections) of $\mathcal{H}_{\mathbb{R}/\mathbb{Z}}$ over this category. In addition, it looks like there are traces of higher homology theory, as the preparations for the proof of Theorem 1 suggest.

### v3

@@ Abstract
The paper contributes to building algebraic foundations of self-organized criticality answering a previously unsolved question about the limiting structure of the extended sandpile group as well as relating it to another limit at the level of classical sandpile groups with respect to certain monomorphisms, and puts forward a concept of canonical sandpile epimorphisms, drawing unexpected consequences about the divisibility properties of the numbers of spanning trees for grid foldings.
@@ 2 Sandpile groups and harmonic functions
where $\mathbb{Z}$ takes the fractional part of the extended recurrent state and $\mathbb{R}/\mathbb{Z}$ is the quotient of $\mathbb{R}$ by the additive subgroup $\mathbb{Z}$ which topologically is a circle.
@@
$0\rightarrow\mathcal{H}_{\mathbb{Z}}(\Gamma)\rightarrow\mathcal{H}_{\mathbb{R}}(\Gamma)\rightarrow\mathcal{H}_{\mathbb{R}/\mathbb{Z}}(\Gamma)\rightarrow\operatorname{Coker}\Delta^{\circ}\rightarrow 0.$
@@
$0\rightarrow\mathcal{H}_{\mathbb{Z}}(\Gamma)\rightarrow\mathbb{Z}^{\partial\Gamma}\rightarrow G(\Gamma)\rightarrow\operatorname{Coker}\Delta^{\circ}\rightarrow 0.$
@@
A diamond is an intersection of $\mathbb{Z}^{2}$ with a rotated by $45$ degrees rectangle $a,b]\times[c,d]\subset\mathbb{R}^{2}$ for some real $a,b,c,d$. Diamonds, among finite convex discrete domains with non-empty interior on the square lattice, are characterized by the property that each vertex of their boundary has more than one neighbor in the complement of the domain. Note that an intersection of two diamonds is again a diamond.
A diamond hull $DH(\Gamma)$ of a finite $\Gamma\subset\mathbb{Z}^{2}$ is the minimal by inclusion diamond containing $\Gamma.$ As the first step, we observe that for convex $\Gamma,$ the restriction homomorphism $\mathcal{H}_{A}(DH(\Gamma))\rightarrow\mathcal{H}_{A}(\Gamma)$ is an isomorphism. This follows from the fact that convexity of $\Gamma$ implies that there is no vertex $v$ in the complement of $\Gamma$ such that it is adjacent to two distinct vertices $w_{1},w_{2}\in\partial\Gamma$ and both $w_{1}$ and $w_{2}$ have only $v$ as an adjacent vertex in the complement. Thus, if $\Gamma$ is not a diamond and connected, it has a vertex $w$ on the boundary adjacent to a unique vertex $v$ in the complement. Consider a new domain $\Gamma^{\prime}=\Gamma\cup\{v\}$ for which $w$ is not a boundary vertex. Then, for $\phi\in\mathcal{H}_{A}(\Gamma),$ there is a unique $\phi^{\prime}\in\mathcal{H}_{A}(\Gamma^{\prime})$ extending $\phi$ to $v$ by $(\Delta\phi)(w),$ i.e. the restriction map $\mathcal{H}_{A}(\Gamma^{\prime})\rightarrow\mathcal{H}_{A}(\Gamma)$ is an isomorphism. Note also that $DH(\Gamma)=DH(\Gamma^{\prime})$ Now, we may consider a finite chain of inclusions
$\Gamma\subset\Gamma^{\prime}\subset\Gamma^{\prime\prime}\subset\dots\subset\Gamma^{(n)}\subset\dots\subset DH(\Gamma),$
where $\Gamma^{(n+1)}$ is obtained by adding one vertex to $\Gamma^{(n)}$ in the same way as we constructed $\Gamma^{\prime}$ from $\Gamma,$ thus each inclusion induces and isomorphism at the level of $A$-valued harmonic functions.
Now we may replace $\Gamma_{1}$ and $\Gamma_{2}$ with their diamond hulls $D_{1}$ and $D_{2}$. Assuming that the inclusion $D_{1}\subset D_{2}$ is strict, there is a vertex $v\in D_{2}\backslash D_{1}$ adjacent to a vertex $w$ in $\partial D_{1}.$ Consider a domain $D_{1}^{\prime}=D_{1}\cup\{v\},$ note that $\partial D_{1}^{\prime}=\partial D_{1}\cup\{v\}.$ Therefore, since we are not requiring the vanishing of the Laplacian on the boundary, an element $\phi\in\mathcal{H}_{A}(D_{1})$ can be extended arbitrarily to $v,$ i.e. the restriction homomorphism $\mathcal{H}_{A}(D_{1}^{\prime})\rightarrow\mathcal{H}_{A}(D_{1})$ is surjective (but not injective for $A\neq 0$). By the previous paragraph, we may replace $D_{1}^{\prime}$ by its diamond hull, resulting in a strictly bigger diamond still contained in $D_{2}.$ Thus, we decompose the restriction homomorphism from $D_{2}$ to $D_{1}$ into a finite chain of restrictions, each of which is surjective. ∎
@@
For a finite $\Gamma\subset\mathbb{Z}^{2}$ the sandpile group $G(\Gamma)$ is generated by adding sand at the boundary.
@@
Remark 2.
For abstract graphs, which are not subgraphs of the lattice, the above statement is not necessarily true. One of the simplest examples is provided by a graph with three non-sink vertices $v_{-},$ $v_{0}$ and $v_{+}$ such that each of the two are connected by a single edge and only $v_{0}$ is connected to the sink (by a single edge), thus $v_{0}$ is the only boundary vertex of this graph. The number of spanning trees on it is $3,$ therefore the sandpile group is isomorphic to $\mathbb{Z}/3\mathbb{Z},$ which has a single non-trivial automorphism given by multiplication by 1,-1, realized by the graph automorphism permuting $v_{-}$ and $v_{+}.$ The neutral element of this sandpile group is invariant under the permutation, and if we add a single grain at the boundary vertex $v_{0},$ the state will remain symmetric after a relaxation, i.e. its result is again the neutral element. We conclude that the subgroup of the sandpile group generated by adding sand at the boundary is trivial in this example.
@@
For an arbitrary finite $\Gamma,$ there is a natural homomorphism from $\mathcal{H}_{\mathbb{R}}(\Gamma)$ to $\widetilde{G}(\Gamma)$ given by taking $\Delta$ of a real-valued harmonic function, which results in a real-valued function supported at the boundary of $\Gamma,$ and taking it modulo the image of $\widetilde{\Delta}$ described at the beginning of this section. Considering the colimit for $\Gamma\rightarrow\mathbb{Z}^{2},$ we still have a homomorphism from $\mathcal{H}_{\mathbb{R}}(\mathbb{Z}^{2})$ to $\lim_{\mathbb{Z}^{2}\leftarrow\Gamma}\widetilde{G}(\Gamma).$ It is easy to see that its kernel is $\mathcal{H}_{\mathbb{Z}}(\mathbb{Z}^{2})$ (this may be done by applying the interpretation of the extended sandpile group as the cockernel of $\widetilde{\Delta}$ and the injectivity of $\Delta$ on finite graphs),i.e. we have an exact sequence
$0\rightarrow\mathcal{H}_{\mathbb{Z}}(\mathbb{Z}^{2})\rightarrow\mathcal{H}_{\mathbb{R}}(\mathbb{Z}^{2})\rightarrow\displaystyle\lim_{\mathbb{Z}^{2}\leftarrow\Gamma}\widetilde{G}(\Gamma)\rightarrow X\rightarrow 0.$
In [6], it was asked: is it true that $X=0$?
@@
Consider an ascending chain of convex discrete domains $\Gamma_{n}$} converging to $\mathbb{Z}^{2}.$ For the fifth and last time we will apply the snake lemma starting with the following diagram, where the rows are short exact sequences (this follows from applying Lemma 3 and then Proposition 1 followed by Lemma 1), and the two squares are commutative (which is the compatibility of $\Psi_{A}$ with the inclusion and projection at the level of coefficients):
[…]
Here, each downward arrow $\Psi_{A},$ for $A=\mathbb{Z}$, $\mathbb{R}$ or $\mathbb{R}/\mathbb{Z},$ stands for the homomorphism given by a formula analogous to ($\star$4) with $\pi_{n}$ being replaced by $\pi^{A}_{n},$ the restriction homomorphism $\mathcal{H}_{A}(\Gamma_{n+1})\rightarrow\mathcal{H}_{A}(\Gamma_{n}).$ The snake lemma for the above diagram gives an exact sequence
$\operatorname{Ker}\Psi_{\mathbb{Z}}\rightarrow\operatorname{Ker}\Psi_{\mathbb{R}}\rightarrow\operatorname{Ker}\Psi_{\mathbb{R}/\mathbb{Z}}\rightarrow\operatorname{Coker}\Psi_{\mathbb{Z}}\rightarrow\operatorname{Coker}\Psi_{\mathbb{R}}\rightarrow\operatorname{Coker}\Psi_{\mathbb{R}/\mathbb{Z}}.$
Note that $\operatorname{Ker}\Psi_{A}$ is by definition the colimit of $\mathcal{H}_{A}(\Gamma_{n})$ which is trivially equal to $\mathcal{H}_{A}(\mathbb{Z}^{2}).$ By Lemma 2, all the homomorphisms $\pi_{n}^{A}$ are surjective, thus we may apply Lemma $\star$4, i.e. all the cokernels in the above six-term sequence vanish, which results in $X=0.$ In other words, $\mathcal{H}_{\mathbb{R}}(\mathbb{Z}^{2})/\mathcal{H}_{\mathbb{Z}}(\mathbb{Z}^{2})$ and $\lim_{\mathbb{Z}^{2}\leftarrow\Gamma}\widetilde{G}(\Gamma)$ are naturally isomorphic. ∎
@@ 3 Monomorphisms and epimorphisms
where $\lfloor\cdot\rfloor$ denotes the integer part, $\phi$ is extended by 0 outside $\Gamma(p,q),$ and $\mu_{p,k}\colon[0,p]\rightarrow[kp,(k+1)p]$ is given by
@@
of the epimorphism and the monomorphism is the identity homomorphism multiplied by $mn$. There are two extreme cases: when $mn$ is coprime with $g(p,q)=|G(\Gamma(p,q))|,$ in which case $G(\Gamma(p,q))$ is realized as a direct summand of $G(\Gamma(mp,nq)),$ or when $mn$ is divisible by $g(p,q),$ and therefore we realize the group $G(\Gamma(p,q))$ as a subgroup of the kernel of the epimorphism, which implies that $(g(p,q))^{2}$ divides $g(mp,nq).$
@@
Table 1: Orders of sandpile groups on $(0,m)\times(0,n)\cap\mathbb{Z}^{2}.$ Observe the divisibility properties in accordance with Theorem 2. For instance, $g(2,4)$ is divisible by $g(2,2),$ but not by $(g(2,2))^{2}$ since the scaling factor $2=1\times 2$ is not 0 modulo $g(2,2)=4;$ on the other hand $(g(2,2))^{2}$ divides $g(4,4).$
A simpler divisibility property which is implied by the existence of epi-/mono-morphism is that $g(p,q)$ always divides $g(mp,nq).$ As it was pointed out to the author, this fact was observed before. The only reference, however, appears to be a webpage [8] in the Online Encyclopedia of Integer Sequences (OEIS), where one may find a contribution by Peter Bala made in 2014, claiming that this property follows from some explicit formula.
The relationship between two rectangles $\Gamma(p,q)$ and $\Gamma(mp,nq)$ represented by a continuous folding map $\Phi_{p,q,m,n}:[0,mp]\times[0,nq]\rightarrow[0,p]\times[0,q],$ which allows to define the corresponding morphisms between their sandpile groups, is a very particular case of a much more general situation. Intuitively, one should imagine that an $mp\times nq$ rectangular piece of grid paper is folded to a $p\times q$ rectangular piece. In this process, every fold is made across a horizontal or vertical line passing through the nodes of the lattice, and the graph structure of the square lattice is preserved. If we cut the bigger piece along the folds, we would obtain $mn$ copies of the $p\times q$ rectangular piece, each of which is placed isometrically on $(0,p)\times(0,q])$ either preserving or reversing the orientation. The defined above sandpile morphisms correspond to a matrix with rows and columns indexed by $\Gamma(p,q)$ and $\Gamma(mp,nq),$ with entries being 0, 11 or 1,-1, with 0 corresponding to either a point of $\Gamma(mp,nq)$ being on a fold (and any point of $\Gamma(p,q)$), or two points $v\in\Gamma(p,q)$ and $w\in\Gamma(mp,nq)$ such that $v\neq\Phi_{p,q,m,n}(w)$, and $\pm 1$ corresponding to the case of $v=\Phi_{p,q,m,n}(w)$ with the sign depending on whether the orientation is preserved or reversed.
In fact, one is allowed to perform the folds not only along the vertical and horizontal lines, but also along the lines of slope $\pm 1.$ This type of lines is distinguished among all other lines since the corresponding reflections preserve the graph structure of $\mathbb{Z}^{2}.$ With this in mind, we give a formal (and, perhaps, not optimal) definition of a grid folding with an adjective “grid” referring to grid paper, some of its instances are shown on Figure 1. This concept is a vast generalization of a $DC$-tiling introduced previously in [7].
Let $\Omega,\Omega^{\prime}\subset\mathbb{R}^{2}$ be two bounded connected domains such that each is equal to the closure of its interior, and let $\Phi\colon\Omega^{\prime}\rightarrow\Omega$ be a continuous map. We call $\Phi$ a grid folding if the following conditions are satisfied:
  • the preimage of the interior of $\Omega$ has $d_{\Phi}$ connected components such that the restriction of $\Phi$ to each is an isometry preserving the lattice;
  • the intersection of the preimage of the boundary of $\Omega$ with the interior of $\Omega^{\prime}$ is a union of segments each of which extends to a line passing through some vertex of the lattice and having a slope equal to either $0,\infty$ or $\pm 1;$
  • for a pair of connected components $\Omega_{1}$ and $\Omega_{2}$ in $\Phi^{-1}(\Omega^{\circ})$ such that the intersection of their closures is a set with more than one point, this intersection is contained in one of the above mentioned lines and the map $(\Phi|_{\Omega_{2}})^{-1}\circ\Phi|_{\Omega_{1}}$ extends to the reflection with respect to this line.
We call $d_{\Phi}$ the degree of the folding $\Phi.$ For instance, for the case of the folding $\Phi_{p,q,m,n}:[0,mp]\times[0,nq]\rightarrow[0,p]\times[0,q],$ its degree is $mn.$ To a folding $\Phi$ one associates the matrix $M_{\Phi}$ indexed by $\Gamma=\Omega^{\circ}\cap\mathbb{Z}^{2}$ and $\Gamma^{\prime}=(\Omega^{\prime})^{\circ}\cap\mathbb{Z}^{2},$ with entries being 0 and $\pm 1,$ depending on wether a point is on the fold, and whether the orientation is preserved or reversed, otherwise. This matrix and its transpose induce a monomorphism and an epimorphism between $G(\Gamma)$ and $G(\Gamma^{\prime}).$ Their composition is equal to the multiplication by $d_{\Phi}$ homomorphism (this is analogous to the composition of cohomology pullback and push-forward associated with a topological covering map). This homomorphism is zero if the degree is divisible by the cardinality of $G(\Gamma)$, which gives the second assertion of the following theorem.
@@
Let $\Gamma$ be a degree $d$ grid folding of $\Gamma^{\prime}.$ Then, the number of spanning trees $g(\Gamma)$ on $\Gamma$ divides the number of spanning trees $g(\Gamma^{\prime})$ on $\Gamma^{\prime}.$ Moreover, if $d$ is divisible by $g(\Gamma)$, then $(g(\Gamma))^{2}$ divides $g(\Gamma^{\prime}).$
@@
For two domains $\Gamma=\Omega^{\circ}\cap\mathbb{Z}^{2}$ and $\Gamma^{\prime}=(\Omega^{\prime})^{\circ}\cap\mathbb{Z}^{2}$ related by a grid folding $\Phi:\Omega^{\prime}\rightarrow\Omega$ such that $\Omega\subset\Omega^{\prime}$ and $\Phi$ restricted to $\Omega$ is the identity map, there is the following commutative diagram:
$G(\Gamma)$$G(\Gamma^{\prime})$$\widetilde{G}(\Gamma)$$\widetilde{G}(\Gamma^{\prime}).$}
@@
It is instructive to see what is the situtation in the one-dimensional case, which is as a general rule dramatically simpler. The extended sandpile group of an integer interval $(0,n)\cap\mathbb{Z}$ for $n\geq 3$ is a two-dimensional torus made of circle-valued arithmetic progressions, and the usual sandpile group is a cyclic group of order $n,$ corresponding to such progressions which take $0\in\mathbb{R}/\mathbb{Z}$ at 0 and $n.$ The direct limit of this groups is $\mathbb{Q}/\mathbb{Z},$ a rational circle corresponding to periodic circle-valued arithmetic progressions with value 0 at the origin, and the inverse limit under the epimorphisms is $\widehat{\mathbb{Z}}=\prod_{l}\mathbb{Z}_{l},$ a direct product of all additive groups of $l$-adic integers, which indeed has no torsion.
Figure 1: Examples of grid foldings of degrees 3, 9 and 2 respectively.
@@ 4 Discussion
There is a striking and not fully explored relation [9] between the sandpile model on a rectangular domain and domino tilings on it. The group playing the role in this relation is not the whole sandpile group of the rectangle, but rather a subgroup of elements symmetric with respect to horizontal and vertical reflections. It seems that with a bit of extra work our Theorem 2 for rectangular foldings may be adapted to the case of such subgroups, and therefore give the analogous divisibility result for the numbers of domino tilings.
The category $\mathcal{D}\mkern-12.0mu\mathcal{D}$ which was passingly introduced in Section 2 imitates by its essence a category of open subsets of some topological space. It makes sense to extend its class of morphisms, leaving the objects intact, by allowing all maps between graphs whose pullbacks on integer-valued functions commute with $\Delta^{\circ}.$ Morally, this allows for partial coverings, thus we may denote such a category as $\mathcal{D}\mkern-12.0mu\mathcal{D}^{et}$ with $et$ standing for “étale”, and $\mathcal{H}_{A}$ is still some kind of a sheaf. It is conceivable that it would be fruitful to study “the universal extended sandpile group” defined as the inverse limit (i.e. global sections) of $\mathcal{H}_{\mathbb{R}/\mathbb{Z}}$ over this category. In addition, it looks like there are traces of higher homology theory, as the preparations for the proof of Theorem 1 suggest.
@@
Acknowledgements
Supported by the Simons Foundation, grant SFI-MPS-T-Institutes-00007697, and the Ministry of Education and Science of the Republic of Bulgaria, grant DO1-239/10.12.2024. The author is grateful to the anonymous referee for the comments which helped to improve the article, and to Andrei Zabolotskii who has pointed out that the first assertion of the original formulation of Theorem 2 was known which motivated a much stronger rephrasing of this theorem not only for rectangles but for arbitrary grid foldings.